# Transfer learning with generative models for object detection on limited datasets

### v1

@@ Abstract
The availability of data is limited in some fields, especially for object detection tasks, where it is necessary to have correctly labeled bounding boxes around each object. A notable example of such data scarcity is found in the domain of marine biology, where it is useful to develop methods to automatically detect submarine species for environmental monitoring. To address this data limitation, the state-of-the-art machine learning strategies employ two main approaches. The first involves pretraining models on existing datasets before generalizing to the specific domain of interest. The second strategy is to create synthetic datasets specifically tailored to the target domain using methods like copy-paste techniques or ad-hoc simulators. The first strategy often faces a significant domain shift, while the second demands custom solutions crafted for the specific task. In response to these challenges, here we propose a transfer learning framework that is valid for a generic scenario. In this framework, generated images help to improve the performances of an object detector in a few-real data regime. This is achieved through a diffusion-based generative model that was pretrained on large generic datasets, and is not trained on the task-specific domain. We validate our approach on object detection tasks, specifically focusing on fishes in an underwater environment, and on the more common domain of cars in an urban setting. Our method achieves detection performance comparable to models trained on thousands of images, using only a few hundreds of input data. Our results pave the way for new generative AI-based protocols for machine learning applications in various domains, for instance ranging from geophysics to biology and medicine.
@@ 1 Introduction
In this paper, we investigate a novel strategy that merges the fields of generative models and object detection. While generative models have been proven successful in generating data for training classifiers [3], their application in the realm of object detection has been less explored. The challenge arises when attempting to generate images suitable for object detectors, as it necessitates the representation of targets within specific layouts, adhering to bounding box positions crucial for supervised training. This issue is formulated as Layout-to-Image (L2I) generation. Traditional generative models, however, faced limitations in generating images with such constraints. To avoid this challenge, some prior approaches employed GANs for performing style transfer, a form of domain adaptation, on the training set [62, 29]. Nonetheless, this method relies on an existing training set, whether synthetic or real, and does not truly generate entirely new images. The first attempt to address the L2I problem was made by Layout2Im [61], but it is only in recent times, notably with GLIGEN [36], that the generation of constrained images has reached a satisfactory level of quality. It is worth emphasizing that the existing literature has explored the problem within the realm of generative models, without application to the object detection task.
@@
We propose to leverage a L2I model to generate abundant labeled data for free, ready for the training of an object detector. To further alleviate the dependence on real labeled data, we intentionally choose not to train the generation model specifically for the real target domain. In particular, we exploit GLIGEN, which is peculiarly built upon existing models pretrained on large-scale web data. This endows GLIGEN with the inherent knowledge derived from other extensive and diverse datasets, facilitating the representation of novel concepts not encountered during its training, aligning seamlessly with our objectives. This sets GLIGEN apart from prior L2I generative models, which were typically confined to be trained on datasets tailored to specific tasks. Furthermore, the inherent value of this strategy lies in its potential applicability to diverse domains without the explicit requirement of training the generator on each individual domain.
In detail, our architecture, illustrated in Fig. 1, adopts a two-step process. Initially, the object detector undergoes training exclusively on generated images, followed by a subsequent fine-tuning stage using only the limited real dataset. The rationale behind this process is to facilitate transfer learning, enabling the model to exploit a large amount of available data in a source domain, here the domain of generated images, to correctly detect data from the target domain, here the domain of real images, for which the amount of training data is limited. Moreover, we designed an optional filtering strategy to assess the impact of the generated dataset on the object detector and to select the most suitable images for its pretraining. To assess the effectiveness of our approach, we evaluate it on two different detection domains: cars and fishes in urban and underwater environments, respectively. The reason for this choice is to compare the effectiveness of the proposed approach on a common domain with high data availability (the former) with another domain with data scarcity and interesting for the scientific community (the latter).
@@ 2 Generative model
Our pipeline is designed to be sufficiently generic, provided that the generator is of the L2I type, and it is capable to generalize to diverse domains. At the best of our knowledge, only GLIGEN satisfies these requirements in the current state-of-the-art.
In the context of GLIGEN, the generation process is guided by an instruction $\textbf{y}=(\textbf{c},\textbf{e})$ that the model follows to create a scene with the desired concepts in specific locations. The instruction comprises a prompt c, a written description of the image content, and a grounding instruction $\textbf{e}=\left[(e_{0},\textbf{l}_{0}),...,(e_{N},\textbf{l}_{N})\right]$. Each $\textbf{l}_{i}$ represents the bounding box coordinates around the object $e_{i}$ that should be depicted within. These objects, referred to as grounding entities, can be described either by text or by a reference real image. In Sec. 6, we will present an ablation study comparing the description of grounding entities through text versus real images. This comparison aims to assess which grounding entity type produces targets best suited for detection in the domain of interest. Regardless of their nature, both kinds of grounding entities are encoded by a pretrained CLIP [47], which is a model trained on vast web-scale data to predict associations between images and corresponding texts, and vice versa. Renowned for its ability to make predictions on concepts not explicitly encountered during training, CLIP encoding is leveraged by GLIGEN, achieving the generalized representation abilities we are seeking. Additionally, associating a reference image as $e_{i}$ with an image edge as the corresponding $\textbf{l}_{i}$ facilitates a consistent style transfer from the real to the generated domain. We leverage this feature to achieve backgrounds that closely resemble the originals.
@@ 3 Object detector
In practice, Faster R-CNN is often trained in a transfer learning setting. One common approach involves pretraining the entire model on a standard detection dataset (e.g., COCO [37]) and subsequently fine-tuning it on the specific domain of interest. However, to avoid biases, we choose not to pretrain on other detection datasets, in order to focus on the contribution of generated data only. Besides, in [28] it is demonstrated that training the entire Faster R-CNN on synthetic images leads to significantly poorer performance compared to training on real data. Conversely, freezing a pretrained feature extractor and exclusively training the remaining parts of the detector results in a considerable performance improvement. Therefore, we freeze a pretrained backbone, and we randomly initialize the RPN weights. Finally, it is worth noting that the backbone can be pretrained on a dataset not necessarily tailored for object detection, thus reducing the need for costly labeling. In detail, in our analyses, we opted to use as backbone a ResNet-50 [26] —a renowned 50 layers deep CNN—pretrained on ImageNet [49], a famous image classification dataset which has become the de facto standard for the pretraining of a CNN [30].
Overall, our Faster R-CNN has 41 370 9114137091141\,370\,91141 370 911 weights, of which 23 454 9122345491223\,454\,91223 454 912 belong to the backbone and are frozen. The remaining weights are trained with Stochastic Gradient Descent (SGD) with momentum, which updates the weights $\theta_{t}$ at time $t$, aiming to minimize a loss function $L$ with the iterative equations:
@@
$b_{t}$ $=\mu b_{t-1}+\nabla_{\theta}L(\theta_{t-1})+\lambda\theta_{t-1}\ ,$
@@
involving the so-called momentum $\mu$, weight decay $\lambda$, learning rate $\gamma$, and $\nabla_{\theta}L$ gradient of $L$. The loss to minimize is formed by the contribution of different terms regarding the positioning of the predicted boxes, the objectness classification and the multiclass classification of each box. In our experiments, we set the value of $\mu=0.9$, while $\lambda$ is $0.001$ during the model pretraining on generated data, and $0.01$ for the fine-tuning on the real data. The reason for the latter is that when the model is trained on a limited dataset size, it is more prone to overfitting. We use higher values of $\mu$, that provide higher regularization, i.e. a penalization to reduce the model complexity.
Regarding the learning rate, we adopted two different scheduling strategies. During the pretraining step, we retain $15\%$ of the generated data as a validation set, a portion of the dataset used to monitor the loss value after each SGD phase over all the remaining training set, i.e. a training epoch. Initially, $\gamma$ is set equal to $0.001$ and if the value of the validation loss do not improve for 5 epochs, $\gamma$ is reduced by a factor of $0.1$. Moreover, we adopt also an early stopping strategy where the detector is trained for a maximum of 200 epochs until the validation loss stops decreasing for at least 10 epochs. In the end, we keep the weights of the model at the epoch with the minor validation loss value. Conversely, during the fine-tuning step, we adopted the common 1x scheduler which has a total fixed number of 12 epochs, and $\gamma$ is initialized with a starting value of $0.001$, before reducing it by a factor $0.1$ at epoch 7 and again at epoch 10. This scheduler does not need a validation set to compute the loss on, which allow us to save data repurposing the real validation set as the test set. We opted to maintain consistent scheduling conditions—and not to further optimize the fine-tuning—to achieve a fair comparison, focusing only on the impact of the pretraining stage in the pipeline.
@@ 4.1 FID score
In detail, features are extracted by an Inception-V3 model [54] pretrained on ImageNet.
@@ 4.2 Precision-Recall
4.3 Mean Average Precision
As we mentioned before, the mAP is the principal metric of interest in the context of object detection. In our analyses, we calculate the mAP on the test set of our real datasets.
@@ 4.3 Mean Average Precision
where $j=1,\dots,M_{i}$ are the ordered predicted boxes for class $i$, $rec_{i,j}$ denotes the recall value calculated for class $i$ on the $j$ predicted box, and $prec(rec_{i,j})$ is the precision value corresponding to $rec_{i,j}$. In other words, the $AP_{i}$ can be seen as the area under an interpolation of the $i$-th precision-recall curve. At each point, the interpolation considers the maximum precision value for which the recall is greater than or equal to the recall at that specific point on the curve. Finally, the mAP is the mean of all $AP_{i}$ over the number of classes $N$. The mAP ranges between 0 and 1: values close to 1 mean that the precision is high also when the recall is increasing. In this case, the number of both false positives and false negatives is low.
@@ 6.1 NuImages—reducing the need for real data
Figure 4: Results of the object detection mAP evaluated on the same NuImages test set, illustrating the impact of pretraining with varying numbers of generated images (x-axis) and subsequent training with different quantities of real images from the NuImages training set (curve color and marker). Solid lines represent results with pretraining on filtered images, while dashed lines depict pretraining on unfiltered images.
@@
In Fig. 4, the x-axis represents the number of generated images used for the initial pretraining, while different colors and markers indicates the numbers of real images used for the fine-tuning step. The y-axis shows the corresponding mAP performances on the real test set for each configuration. Besides, to evaluate the usefulness of our filter, we followed two distinct approaches: one involving the selection of generated images through a Precision-Recall filter (solid curves), and the other employing all generated images, regardless of their quality (dashed curves). Specifically, for both precision and recall, we set the filtering thresholds to 1, aiming to preserve only the most accurate representations of the ground truth during the generation process (this process discarded around 56% of all the images).
@@
In another test, we also evaluate the influence of the FID score on the training of the object detector. We arranged the unfiltered generated images into three mutually exclusive subsets with varying FID scores: 47, 105, and 148. Each subset consisted of 3 000 images for pretraining, with an additional 530 images used for validation. All these configurations are then involved in a fine-tuning step on the same 300 real images of Fig. 4. At the end, we found that there are no significant improvements in training the detector on datasets with different FID scores. Specifically, for all three cases we obtained really similar test detection mAPs comparable with the results of Fig. 4. This is an evidence that, even if the FID score serves as a valuable metric for assessing the overall similarity between real and generated distributions, it may not be the most suitable metric to measure the impact of generated data in object detection tasks. Specifically, the FID calculates distribution distances considering the images as a whole, and not particularly regarding the quality of each single target. For example, we can hypothetically consider two distinct detection datasets having exactly identical images, but one with the correct labeling and the other one with a completely erroneous one. In this scenario, the FID distance between the two datasets would be 0, but their impact on the training of an object detector would be completely different, catastrophic for the wrongly labelled dataset.
6.2 OzFish—ablation study on different pretraining datasets
@@
Figure 5: Object detection mAP results on the OzFish real test set, for models pretrained on different quantities of unfiltered generated images (x-axis) and fine-tuned on varying numbers of OzFish training images (specified by the curve color and marker). Dotted curves indicate models pretrained on synthetic copy-paste images. Dashed and solid lines represent the use of images and text as grounding entities, respectively. The black constant solid line reports for reference the mAP value of a model pretrained on COCO and fine-tuned on all the 1 500 OzFish training images.
In this section we discuss the application of our pipeline to the domain of OzFish. We present a comparison between our method and other prominent state-of-the-art approaches. Moreover, in this domain we decided not to focus on the effects of the filter because we observed a fast convergence already using the unfiltered images.
@@
We employed the three datasets outlined in Sec. 5 and exemplified in Fig. 3. Similar to the generated dataset used for the NuImages scenario, we maintained an 85% of the generated images for pretraining and reserved the remaining 15% for validation. As in Fig. 4, in Fig. 5 we plot the mAP trend for an increasing number of pretraining generated images (x-axis) and fine-tuning real images (curve color). For all cases, the mAP is computed on the same real test set of 265 OzFish images.
@@
The first insightful observation arises when comparing various image generation strategies. We observe that each method’s curve appears completely above the previous ones. Specifically, the copy-paste approach (dotted curve) exhibits the highest performance, followed by the cases where grounding entities are images (dashed) and text (solid). This is probably because the copy-paste dataset incorporates true fishes features, which may prove more effective than the generated ones. However, it is crucial to note that the curves appear to be sorted by an increasing level of their generation complexity. It is important to recall that the copy-paste approach synthesizes images by relying on mask segmentation, a process that is inherently resource-intensive, even more than labeling a dataset for object detection. In contrast, our methods stand out for their simplicity, requiring only target pictures for image grounding or even simpler text descriptions. For instance, under the configuration with 9 000 pretraining images and 300 real images, the mAP gap between text grounding and copy-paste methods is just 0.05, which is quite tolerable, considering the substantial reduction in the burden on human intervention.
@@
Finally, in Fig. 5 we also establish a reference configuration whose weights are initialized based on the COCO dataset as in [48]. This allows to benchmark the performances of our generated datasets against the most prominent state-of-the-art dataset for the pretraining of any object detector. In detail, the reference model is pretrained on ImageNet and fine-tuned on COCO. We additionally fine-tune it on the whole OzFish dataset, keeping also in this case the backbone frozen as detailed in Sec. 3. Our model surprisingly approaches the performance of the COCO-pretrained reference (black line in Fig. 5) with only 300 OzFish examples (red curve), reducing the number of real labeled images by almost three orders of magnitude: from 118K COCO images plus 1 500 OzFish images to only 300 OzFish images supported by 9 000 generated ones. Besides, with this specific training setting, when using the entire dataset of 1 500 real images, we manage to surpass the reference with only 1 500 generated images (green curve). These are insights that the generated data is better suited for the transfer learning task in this context, compared to the general-purpose COCO dataset. This is encouraging for the widespread adoption of generative models in object detection tasks.
@@
Figure 6: Side-by-side illustrations comparing examples of object detector predictions (in red) with the corresponding testing ground truth (in green) under our two scenarios. In the left column (a, c), Faster R-CNN models are initialized with ImageNet and fine-tuned using a minimal dataset of 300 real training images. On the right, models undergo a preliminary pretraining on 9 000 generated images (from the image grounding dataset in the (d) case), followed by a fine-tuning on the same 300 real images. We can observe that our approach shows increased performances in finding more targets and refining redundant detection boxes, even in rainy or turbid environments.

### v2

@@ Abstract
The availability of data is limited in some fields, especially for object detection tasks, where it is necessary to have correctly labeled bounding boxes around each object. A notable example of such data scarcity is found in the domain of marine biology, where it is useful to develop methods to automatically detect submarine species for environmental monitoring. To address this data limitation, the state-of-the-art machine learning strategies employ two main approaches. The first involves pretraining models on existing datasets before generalizing to the specific domain of interest. The second strategy is to create synthetic datasets specifically tailored to the target domain using methods like copy-paste techniques or ad-hoc simulators. The first strategy often faces a significant domain shift, while the second demands custom solutions crafted for the specific task. In response to these challenges, here we propose a transfer learning framework that is valid for a generic scenario. In this framework, generated images help to improve the performances of an object detector in a few-real data regime. This is achieved through a diffusion-based generative model that was pretrained on large generic datasets. With respect to the state-of-the-art, we find that it is not necessary to fine tune the generative model on the specific domain of interest. We believe that this is an important advance because it mitigates the labor-intensive task of manual labeling the images in object detection tasks. We validate our approach focusing on fishes in an underwater environment, and on the more common domain of cars in an urban setting. Our method achieves detection performance comparable to models trained on thousands of images, using only a few hundreds of input data. Our results pave the way for new generative AI-based protocols for machine learning applications in various domains, for instance ranging from geophysics to biology and medicine.
@@ 1 Introduction
In this paper, we investigate a novel strategy that merges the fields of generative models and object detection. While generative models have been proven successful in generating data for training classifiers [40], their application in the realm of object detection has been less explored. The challenge arises when attempting to generate images suitable for object detectors, as it necessitates the representation of targets within specific layouts, adhering to bounding box positions crucial for supervised training. This issue is formulated as Layout-to-Image (L2I) generation. Traditional generative models, however, faced limitations in generating images with such constraints. To avoid this challenge, some prior approaches employed GANs for performing style transfer, a form of domain adaptation, on the training set [41, 42]. Nonetheless, this method relies on an existing training set, whether synthetic or real, and does not truly generate entirely new images. The first attempt to address the L2I problem was made by Layout2Im [43], but it is only in recent times, notably with GLIGEN [44], that the generation of constrained images has reached a satisfactory level of quality.
It is worth emphasizing that the existing literature has mostly explored the problem within the realm of generative models, rather than its application to the object detection task. The work [45] also explored the potential of training an object detector using generated images. The authors fine-tuned their generative model on a big real dataset from a very specific domain. However, this requires data with time-consuming bounding-box labeling, potentially leading to a cyclical issue where large, labor-intensive data may still be needed to train the generative model instead of the detector. By contrast, we aim to mitigate this problem, by tackling the data scarcity on generic domains with the aid of pre-trained generative models with L2I capabilities, to generate abundant labeled data for free, ready for the training of an object detector. Indeed, we intentionally choose to use a pretrained generative model to further alleviate the dependence on real labeled data. Thus, an important novelty of our work is the absence of the generative model training on the specific task domain.
@@
In particular, we exploit GLIGEN, which is peculiarly built upon existing models pretrained on large-scale web data. This endows GLIGEN with the inherent knowledge derived from other extensive and diverse datasets, facilitating the representation of novel concepts not encountered during its training, aligning seamlessly with our objectives. This sets GLIGEN apart from prior L2I generative models, which were typically confined to be trained on datasets tailored to specific tasks. Furthermore, the inherent value of this strategy lies in its potential applicability to diverse domains without the explicit requirement of training the generator on each individual domain.
In detail, our architecture, illustrated in Fig. 1, adopts a two-step process. Initially, the object detector undergoes training exclusively on generated images, followed by a subsequent fine-tuning stage using only the limited real dataset. The rationale behind this process is to facilitate transfer learning, enabling the model to exploit a large amount of available data in a source domain, here the domain of generated images, to correctly detect data from the target domain, here the domain of real images, for which the amount of training data is limited. Moreover, we designed an optional filtering strategy to assess the impact of the generated dataset on the object detector and to select the most suitable images for its pretraining. To assess the effectiveness of our approach, we evaluate it on two different detection domains: cars and fishes in urban and underwater environments, respectively. The reason for this choice is to compare the effectiveness of the proposed approach on a common domain with high data availability (the former) with another domain with data scarcity and interesting for the scientific community (the latter). Nevertheless, the domain choice comes without loss of generality, since our framework can in principle be applied to other common or uncommon domains.
@@ 2 Generative model
Our pipeline is designed to be sufficiently generic, provided that the generator is of the L2I type, and it is capable to generalize to diverse domains. At the best of our knowledge, only GLIGEN satisfies these requirements in the current state-of-the-art. This is because GLIGEN leverages CLIP, a model trained on vast web-scale data to predict associations between images and corresponding texts, and vice versa. Renowned for its ability to make predictions on concepts that are not explicitly encountered during training, CLIP encoding is leveraged by GLIGEN, achieving the generalized representation abilities we are seeking.
Specifically, the generation process of GLIGEN is guided by an instruction $\textbf{y}=(\textbf{c},\textbf{e})$ that the model follows to create a scene with the desired concepts in specific locations. The instruction comprises a prompt c, a written description of the image content, and a grounding instruction $\textbf{e}=\left[(e_{0},\textbf{l}_{0}),...,(e_{N},\textbf{l}_{N})\right]$. Each $\textbf{l}_{i}$ represents the bounding box coordinates around the object $e_{i}$ that should be depicted within. These objects, referred to as grounding entities, can be described either by text or by a reference real image. In Sec. 6, we will present an ablation study comparing the description of grounding entities through text versus real images. This comparison aims to assess which grounding entity type produces targets best suited for detection in the domain of interest. Regardless of their nature, both kinds of grounding entities are encoded by the pretrained CLIP [46]. Additionally, associating a reference image as $e_{i}$ with an image edge as the corresponding $\textbf{l}_{i}$ facilitates a consistent style transfer from the real to the generated domain. We leverage this feature to achieve backgrounds that closely resemble the originals.
@@ 3 Object detector
In practice, Faster R-CNN is often trained in a transfer learning setting. One common approach involves pretraining the entire model on a standard detection dataset (e.g., COCO [48]) and subsequently fine-tuning it on the specific domain of interest. However, to avoid biases, we choose not to pretrain on other detection datasets, in order to focus on the contribution of generated data only. Besides, in [49] it is demonstrated that training the entire Faster R-CNN on synthetic images leads to significantly poorer performance compared to training on real data. Conversely, freezing a pretrained feature extractor and exclusively training the remaining parts of the detector results in a considerable performance improvement. Therefore, we freeze a pretrained backbone, and we randomly initialize the RPN weights. Finally, it is worth noting that the backbone can be pretrained on a dataset not necessarily tailored for object detection, thus reducing the need for costly labeling. In detail, in our analyses, we opted to use as backbone a ResNet-50 [50] —a renowned 50 layers deep CNN—pretrained on ImageNet [51], a famous image classification dataset which has become the de facto standard for the pretraining of a CNN [52]. An in-depth description of the ResNet-50 backbone and the Faster R-CNN model are detailed in Appendix A.
Overall, our Faster R-CNN has 41 370 9114137091141\,370\,91141 370 911 weights, of which 23 454 9122345491223\,454\,91223 454 912 belong to the backbone and are frozen. The remaining weights are trained with Stochastic Gradient Descent (SGD) with momentum, which updates the weights $\theta_{t}$ at time $t$, aiming to minimize a loss function $L$ with the iterative equations [53]:
@@
$b_{t}$ $=\mu b_{t-1}+\nabla_{\theta}L(\theta_{t-1})+\lambda\theta_{t-1}\ ,$ (1)
@@
involving the so-called momentum $\mu$, weight decay $\lambda$, learning rate $\gamma$, and $\nabla_{\theta}L$ gradient of $L$. The loss to minimize is formed by the contribution of different terms regarding the positioning of the predicted boxes, the objectness classification and the multiclass classification of each box. In our experiments, we set the value of $\mu=0.9$, while the weight decay $\lambda$ is $0.001$ during the model pretraining on generated data, and $0.01$ for the fine-tuning on the real data. The reason for the latter is that when the model is trained on a limited dataset size, it is more prone to overfitting. We use higher values of $\lambda$, that provide higher regularization, i.e. a penalization to reduce the model complexity.
Regarding the learning rate, we adopted two different scheduling strategies, depending on the availability of images for a validation set—a portion of the dataset used to monitor the loss value after each SGD phase (epoch) over the entire training set. In the fine-tuning stage, due to limited real data, we opted for the common 1x scheduler. This scheduler does not require a validation set, allowing us to maximize training data usage. Conversely, during the pre-training stage, where we can generate as much fake data as needed, we dedicate $15\%$ of the generated data as a validation set. At the beginning of both stages, the learning rate $\gamma$ is initialized with a starting value of $0.001$. Then, the pretraining step scheduler reduces $\gamma$ by a factor of $10$ if the value of the validation loss does not improve for 5 epochs. Moreover, we adopt also an early stopping strategy where the detector is trained for a maximum of 200 epochs until the validation loss stops decreasing for at least 10 epochs. In the end, we keep the weights of the model at the epoch with the minor validation loss value. During the fine-tuning step, the 1x scheduler has a total fixed number of 12 epochs, and $\gamma$ is reduced by a factor of $10$ at epoch 7 and again at epoch 10. The fixed 1x scheduler should not be considered as a limitation, instead, it has been chosen to achieve a fair comparison, focusing only on the impact of the pretraining stage in the pipeline.
@@ 4.1 FID score
In detail, we follow the method introduced by [54] and widely adopted in the later literature, which involves extracting features with an Inception-V3 model [55] pretrained on ImageNet.
@@ 4.2 Precision-Recall
4.3 Mean Average Precision and F-score
As we mentioned before, the mAP is the principal metric to measure accuracy in the context of object detection. In our analyses, we calculate the mAP on the test set of our real datasets.
@@ 4.3 Mean Average Precision and F-score
where $j=1,\dots,M_{i}$ are the ordered predicted boxes for class $i$, $rec_{i,j}$ denotes the recall value calculated for class $i$ on the $j$ predicted box, and $prec(rec_{i,j})$ is the precision value corresponding to $rec_{i,j}$. In other words, the $AP_{i}$ can be seen as the area under an interpolation of the $i$-th precision-recall curve. At each point, the interpolation considers the maximum precision value for which the recall is greater than or equal to the recall at that specific point on the curve. Finally, the mAP is the mean of all $AP_{i}$ over the number of classes $N$. By definition, there is a trade-off between precision and recall: increasing recall will inevitably cause non-relevant results to be detected, decreasing precision. The mAP ranges between 0 and 1: values close to 1 mean that the precision is high also when the recall is increasing. In this case, the number of both false positives and false negatives is low.
Secondly, the performance of our models are evaluated by an F1 score, which considers all the detections having confidence greater than 0.5, and represents both precision and recall in one metric:
$F_{1}$ $=2\,\frac{prec\cdot rec}{prec+rec}$ (5)
@@ 6.1 NuImages—reducing the need for real data
Figure 4: Results of the object detection mAP evaluated on the same NuImages test set, illustrating the impact of pretraining with varying numbers of generated images (x-axis) and subsequent training with different quantities of real images from the NuImages training set (curve color and marker shape). Solid lines with filled markers represent results with pretraining on filtered images, while dashed lines with empty markers depict pretraining on unfiltered images (in the right part of the legend we report only one marker shape for compactness).
@@
In Fig. 4, the x-axis represents the number of generated images used for the initial pretraining, while different colors and marker shapes indicates the numbers of real images used for the fine-tuning step. The y-axis shows the corresponding mAP performances on the real test set for each configuration. Besides, to evaluate the usefulness of our filter, we followed two distinct approaches: one involving the selection of generated images through a Precision-Recall filter (solid curves and filled markers), and the other employing all generated images, regardless of their quality (dashed curves and empty markers). Specifically, for both precision and recall, we set the filtering thresholds to 1, aiming to preserve only the most accurate representations of the ground truth during the generation process (this process discarded around 56% of all the images).
@@
In another test, we also evaluate the influence of the FID score on the training of the object detector. We arranged the unfiltered generated images into three mutually exclusive subsets with varying FID scores: 47, 105, and 148. Each subset consisted of 3 000 images for pretraining, with an additional 530 images used for validation. All these configurations are then involved in a fine-tuning step on the same 300 real images of Fig. 4. At the end, we found that there are no significant improvements in training the detector on datasets with different FID scores. Specifically, for all three cases we obtained really similar test detection mAPs comparable with the results of Fig. 4. This is an evidence that, even if the FID score serves as a valuable metric for assessing the overall similarity between real and generated distributions, it may not be the most suitable metric to measure the impact of generated data in object detection tasks. Specifically, it is important to recognize that FID only assesses the image quality independently from specific downstream tasks like object detection. Indeed, the FID calculates distribution distances considering the images as a whole, rather than focusing on the quality of individual targets, which is the primary concern of the object detector’s feature extractor. For example, we can hypothetically consider two distinct detection datasets having exactly identical images, but one with the correct labeling and the other one with a completely erroneous one. In this scenario, the FID distance between the two datasets would be 0, but their impact on the training of an object detector would be completely different, catastrophic for the wrongly labelled dataset.
6.2 OzFish—comparing with different state-of-the-art techniques and testing universality
@@
Figure 5: Object detection mAP results on the OzFish real test set, for models pretrained on different quantities of unfiltered generated images (x-axis) and fine-tuned on varying numbers of OzFish training images (specified by the curve color). Dotted curves with round markers indicate models pretrained on synthetic copy-paste images. Dot-dashed and dashed lines, with squares and diamond markers, represent the use of images and text as grounding entities, respectively. The cyan constant solid line at around $0.6$ mAP reports for reference the performance of a model pretrained on COCO and fine-tuned on all the 1 500 OzFish training images. The colored solid lines with triangular markers are references that use a standard data augmentation approach.
@@
In this section we discuss the application of our pipeline to the domain of OzFish. We present a comparison between our method and other prominent state-of-the-art approaches. Moreover, in this domain we decided not to focus on the effects of the filter because we observed a fast convergence already using the unfiltered images. We employed the three datasets outlined in Sec. 5 and exemplified in Fig. 3. Similar to the generated dataset used for the NuImages scenario, we maintained an 85% of the generated images for pretraining and reserved the remaining 15% for validation. As in Fig. 4, in Fig. 5 we plot the mAP trend for an increasing number of pretraining generated images (x-axis) and fine-tuning real images (curve color). For all cases, the mAP is computed on the same real test set of 265 OzFish images.
@@
The first insightful observation arises when comparing various image generation strategies. We observe that each method’s curve appears completely above the previous ones. Specifically, the state-of-the-art copy-paste approach (dotted curve with round markers) exhibits the highest performance, followed by the cases where grounding entities are images (dot-dashed curve with square markers) and text (dashed curve with diamond markers). This is probably because the copy-paste dataset incorporates true fishes features, which may prove more effective than the generated ones. However, it is crucial to note that the curves appear to be sorted by an increasing level of their generation complexity. It is important to recall that the copy-paste approach synthesizes images by relying on mask segmentation, a process that is inherently resource-intensive, even more than labeling a dataset for object detection. In contrast, our methods stand out for their simplicity, requiring only target pictures for image grounding or even simpler text descriptions. For instance, under the configuration with 9 000 pretraining images and 300 real images, the mAP gap between text grounding and copy-paste methods is just 0.05, which is quite tolerable, considering the substantial reduction in the burden on human intervention.
@@
We also compare our model with the classical data augmentation techniques illustrated in Sec. 1 (colored solid lines with triangular markers in Fig. 5). We augment independently each of the considered real sub-datasets, and we directly train the object detector (pretrained only on ImageNet) on a mix of real and augmented images, without our 2-step transfer learning. Moreover, we establish another reference configuration whose weights are initialized based on the COCO dataset as in [47]. This allows to benchmark the performances of our generated datasets against the state-of-the-art method to address data scarcity, that is to perform transfer learning using a model pretrained on the most prominent dataset for object detection. In detail, the reference model is pretrained on ImageNet and then on COCO. We additionally fine-tune it on the whole OzFish dataset, keeping also in this case the backbone frozen as detailed in Sec. 3. Our model surprisingly approaches the performance of the COCO-pretrained reference (cyan constant solid line at around $0.6$ mAP in Fig. 5) with only 300 OzFish examples (red curve), reducing the number of real labeled images by almost three orders of magnitude: from 118K COCO images plus 1 500 OzFish images to only 300 OzFish images supported by 9 000 generated ones. Besides, with this specific training setting, when using the entire dataset of 1 500 real images, we manage to surpass the reference with only 1 500 generated images (green curve). These are insights that the generated data is better suited for the transfer learning task in this context, compared to the general-purpose COCO dataset. This is encouraging for the widespread adoption of generative models in object detection tasks.
Furthermore, in Fig. 6, we test the universality of our framework by comparing our approach across different object detectors. We recall that the object detector used in previous evaluations is a Faster R-CNN with a ResNet-50 backbone (labeled as Faster Resnet in the legend). This is compared with two different scenarios. In the first, the same Faster R-CNN model is used, but with a different backbone network: MobileNet_v3 [67] (labeled as Faster Mobile). The second scenario employs the same ResNet-50 backbone, but with a completely different detection model: Fully Convolutional One-Stage object detector (FCOS[68] (labeled as FCOS Resnet). Technical details of these models can be found in the Appendix A. We observe that all the three models exhibit a similar behavior, demonstrating that our approach is independent of the chosen object detector and highlighting the beneficial effect of pretraining on generated images in every scenario. Finally, all the main metrics (mAP, precision, recall and F1 score) for some relevant trainings of the object detector on the OzFish dataset are reported in table 1, along with the results on NuImages.
Figure 6: Comparison of the Faster R-CNN model used in Fig. 5 with other models. All tests use 300 real images with pretraining on text-grounded generated images.
Table 1: Metrics for the two considered datasets, for 50 and 1 50015001\,5001 500 real images and for 0 and 9 00090009\,0009 000 generated images. The confidence threshold to compute precision, recall and F1 is fixed at 0.5. The OzFish generated dataset is the image grounding one.
Dataset Real im. Generated im. Prec. Rec. F1 mAP
NuImages 50 0 0 0 - 0.74e-6
9 00090009\,0009 000 0.51 0.56 0.53 0.55
1 50015001\,5001 500 0 0.35 0.54 0.42 0.43
9 00090009\,0009 000 0.52 0.67 0.59 0.66
OzFish 50 0 0 0 - 0.024
9 00090009\,0009 000 0.48 0.51 0.50 0.48
1 50015001\,5001 500 0 0.41 0.54 0.46 0.46
9 00090009\,0009 000 0.57 0.65 0.61 0.64
@@
Figure 7: Side-by-side illustrations comparing examples of object detector predictions (in red) with the corresponding testing ground truth (in green) under our two scenarios. In the left column (a, c), Faster R-CNN models are initialized with ImageNet and fine-tuned using a minimal dataset of 300 real training images. On the right, models undergo a preliminary pretraining on 9 000 generated images (from the image grounding dataset in the OzFish case), followed by a fine-tuning on the same 300 real images. We can observe that our approach shows increased performances in finding more targets and refining redundant detection boxes, even in rainy or turbid environments.